0.145 gram of hydrocarbon is heated with dry copper oxide and 224 cm cube of CO2 was collected at STP the empirical formula for hydrocarbon 

Grade:11

1 Answers

Arun
5 years ago
Sample X contains carbon and hydrogen only. (ii) Copper (II) oxide acts as oxidising agent.
(iii) 1. 22400 cm3 of CO2 at STP = 1 mole
∴ 224cm3 ofCO2 atSTP= 224 =0.01mole. 22400
2. 22400 cm3 of carbon dioxide contains carbon = 12 g
∴ 224 cm3 of carbon dioxide contains carbon = 12 × 224 22400
  = 0.12g 3. MassofhydrogeninsampleX=Massofhydrocarbon–
∴ Wt. of carbon in sample X = 0.12 g. Mass of carbon = (0.145 – 0.12) g = 0.025 g.
 
Element.      %age wt.         At. wt.         Relative no. of moles.         Simple ratio of atoms
C.                0.120 g.             12.               0.120 ÷ 12 = 0.01.              0.01 ÷ 0.01 = 1 or 2
H.                0.025 g.               1.               0.025 ÷ 1 = 0.025.            0.025 ÷ 0.01 = 2.5 or 5
 
hence empirical formula = C2H5